# Localizing Dynamically-Formed Black Hole Binaries in Milky Way Globular Clusters with LISA

### v1

@@ 4 Discussion
To evaluate the measurement accuracy achievable with LISA, we performed a Fisher matrix analysis on the simulated BBH population, as illustrated in Figures 3 and 4. Particularly, Figure 3 maps the intrinsic parameters, $(a,1-e)$, of the simulated Milky Way GC BBH population and uses the background color to show their different parameter measurement errors (5-year observation); Figure 4 depicts the measurement error of $\{f_{\rm orb},e,M,q,R,2a_{s}\}$ for all the simulated BBHs with $\rm SNR>3$, across all the 10 realizations, for 5-year observation. As illustrated in both figures, most of the BBHs with $\rm SNR>3$ have well-constrained orbital frequency and eccentricity, reaching $\delta f_{\rm orb}/f_{\rm orb}\sim 10^{-7}-10^{-5}$ and $\delta e/e<1$. Furthermore, eccentricity can, in general, enhance the measurement accuracy, with $\delta e/e$ reaching $\sim 10^{-3}-0.1$ for detectable BBHs with $e>0.7$. On the other hand, BBHs with $\rm SNR>20$ can have a total mass measurement accuracy of $\delta M/M\lesssim 10\%$, but most of the marginally detectable sources may not have a well-constrained total mass and mass ratio. Also, we expect BBHs with $\rm SNR>20$ to be confidently localized with an angular resolution of $\sim 10-100$ arcmins, which enables us to localize these sources in specific Milky Way GCs. We further verified the Fisher matrix results against Bayesian analysis for an example system, as shown in Figure 5.
Although our collective sample of GC simulations effectively spans the full parameter space of Galactic GCs (Kremer et al., 2020), this simulation suite is a grid, which inevitably means some specific observed GCs do not necessarily have a strong “one-to-one” model match. For example, Ref. Rui et al. (2021), demonstrated how the CMC Catalog may be augmented with additional models to match particular GCs with observed properties lying in between grid points. In another recent study, Ye et al. (2022) showed that especially massive GCs like 47 Tuc (NGC 104) may require additional adjustments (in particular variations to the initial density profile and initial stellar mass function) to produce a precise model match. The ideal solution is to produce a separate model for every single Galactic cluster (for some examples, see Kremer et al., 2018b, 2019a; Ye et al., 2022, 2024); however, this is computationally expensive, and outside the scope of the current study. We tested the LISA predictions from our best-fit model for NGC 104 with the predictions from the more precise model in Ye et al. (2022), and found that the number of LISA sources is consistent within a factor of $3$ for SNR$>1$.
@@
ZX acknowledges partial support from the Bhaumik Institute for Theoretical Physics summer fellowship. ZX and SN acknowledge the partial support from NASA ATP Grant No. 80NSSC20K0505 and from NSF-AST Grant No. 2206428 and thank Howard and Astrid Preston for their generous support. Further, ZX, KK and SN thank LISA Sprint 2024 for organizing an interactive and productive meeting.

### v2

@@ 4 Discussion
To evaluate the measurement accuracy achievable with LISA, we performed a Fisher matrix analysis on the simulated BBH population, as illustrated in Figures 3 and 4. Particularly, Figure 3 maps the intrinsic parameters, $(a,1-e)$, of the simulated Milky Way GC BBH population and uses the background color to show their different parameter measurement errors (5-year observation); Figure 4 depicts the measurement error of $\{f_{\rm orb},e,M,q,R,2a_{s}\}$ for all the simulated BBHs with $\rm SNR>3$, across all the 10 realizations, for 5-year observation. As illustrated in both figures, most of the BBHs with $\rm SNR>3$ have well-constrained orbital frequency and eccentricity, reaching $\delta f_{\rm orb}/f_{\rm orb}\sim 10^{-7}-10^{-5}$ and $\delta e/e<1$. Furthermore, eccentricity can, in general, enhance the measurement accuracy, with $\delta e/e$ reaching $\sim 10^{-3}-0.1$ for detectable BBHs with $e>0.7$. On the other hand, BBHs with $\rm SNR>20$ can have a total mass measurement accuracy of $\delta M/M\lesssim 10\%$, but most of the marginally detectable sources may not have a well-constrained total mass and mass ratio. Also, we expect BBHs with $\rm SNR>20$ to be confidently localized with an angular resolution of $\sim 10-100$ arcmins, which enables us to localize these sources in specific Milky Way GCs.
[…]
Although our collective sample of GC simulations effectively spans the full parameter space of Galactic GCs (Kremer et al., 2020), this simulation suite is a grid, which inevitably means some specific observed GCs do not necessarily have a strong “one-to-one” model match. For example, Ref. Rui et al. (2021), demonstrated how the CMC Catalog may be augmented with additional models to match particular GCs with observed properties lying in between grid points. In another recent study, Ye et al. (2022) showed that especially massive GCs like 47 Tuc (NGC 104) may require additional adjustments (in particular variations to the initial density profile and initial stellar mass function) to produce a precise model match. The ideal solution is to produce a separate model for every single Galactic cluster (for some examples, see Kremer et al., 2018b, 2019a; Ye et al., 2022, 2024); however, this is computationally expensive, and outside the scope of the current study. We tested the LISA predictions from our best-fit model for NGC 104 with the predictions from the more precise model in Ye et al. (2022), and found that the number of LISA sources is consistent within a factor of $3$ for SNR$>1$. Furthermore, the multimodality of the posterior, particularly in angular parameters, can pose additional challenges for accurate sky localization (see, e.g., Marsat et al., 2021). However, a comprehensive analysis of this effect is beyond the scope of the present work.
@@
The authors thank the anonymous referee for their valuable feedback and helpful suggestions. ZX acknowledges partial support from the Bhaumik Institute for Theoretical Physics summer fellowship. ZX and SN acknowledge the partial support from NASA ATP Grant No. 80NSSC20K0505 and from NSF-AST Grant No. 2206428 and thank Howard and Astrid Preston for their generous support. Further, ZX, KK and SN thank LISA Sprint 2024 for organizing an interactive and productive meeting.
Figure 6: Histogram of Milky Way GC angular diameters and their hosted BBH counts. Here we show the distribution of GC angular diameters (left), taken from the Milky Way Star Clusters Catalog (Kharchenko et al., 2012, 2013; Schmeja et al., 2014; Scholz et al., 2015), and the number of hosted BBH systems with SNR $>1$ in each GC diameter bin (right), based on 10 realizations from the CMC Cluster Catalog. The difference between the left and right panels arises because detectable BBH systems tend to be hosted by more massive and nearby GCs, which also have a larger angular size.
Figure 7: Parameter Measurement error of an example BBH system in Milky Way Globular Clusters. Here we adopt the same method as Figure 5, but consider a different BBH system with $m_{1}=17.6$ ${}_{\odot},$ $m_{2}=16.5$ ${}_{\odot},$ $f_{\rm orb}=1.572852\times 10^{-6}$ Hz, $e=0.9668,$ $R=10$ kpc, $\Phi=329.2^{\circ},\,\Theta=64.5^{\circ},\,\phi=341.4^{\circ},\,\theta=149.8^{\circ},\,\psi=14.8^{\circ}$ ($\rm SNR\sim 9$).
Table 1: Sky location and distance measurement accuracy of representative systems in the mock Milky Way GC catalog, compared with their host cluster size (5-yr observation)
Host Cluster Cluster Size $\delta\Omega_{\rm Bayesian}$ $\delta\Omega_{\rm Fisher}$ $\delta R$ $M$ $a$ $e$ SNR
[deg2] [deg2] [deg2] [kpc] [$\rm M_{\odot}$] [au] (5-yrs)
NGC104 2.84 0.40 0.073 0.25 29.5 0.01102 0.2803 94
NGC5139 2.13 1.00 0.17 2.16 41.9 0.06332 0.8969 77
NGC6341 0.52 1.85 0.38 18.0 21.2 0.00750 0.1671 35
Note. — Here $\delta\Omega_{\rm Bayesian}$ represents the sky area of 68% credible regions of the marginal posterior densities (see, e.g., the bottom right panel of Figure 5); $\delta\Omega_{\rm Fisher}=\pi a_{s}b_{s}$ represents the area of sky error ellipsoid, as estimated by Fisher matrix analysis; $\delta R$ shows the distance measurement accuracy, as of $3\sigma$ level from Bayesian analysis; $M,a,e$ represent the total mass, semi-major axis, and eccentricity of the binary system, respectively. For comparison purposes, we show the name and sky area of each binary’s host cluster (inferred from their angular radius in the Milky Way Star Clusters Catalog, Kharchenko et al., 2012, 2013; Schmeja et al., 2014; Scholz et al., 2015).